if you are dealt two cards successively (with replacement of the first)from a standard 52 card deck, find the probability of getting a face card on the first card and an ace on the second. (face cards include the king, queen and jack)

Answer by stanbon(75887)

if you are dealt two cards successively (with replacement of the first)from a standard 52 card deck, find the probability of getting a face card on the first card and an ace on the second. (face cards include the king, queen and jack)
---------------------------
Replacement results in the selection results being independent.
------
P(face AND ace) = (12/52)(4/52) = (3/13)(1/13) = 4/169
--------------------------------
Cheers,
Stan H.
===============